Compute the value of the discriminant and give the number of real solutions to the quadratic equation.
5x^2 - x - 3 = 0
Please help!!!!!

Answer by mukhopadhyay(490)

5x^2 - x - 3 = 0
Discriminant for ax^2+bx+c=0 is b^2-4ac
Here a=5, b=-1, and c=-3
b^2-4ac = 1+60 = 61
Roots are x=[-b+sqrt(discriminant)]/2a and x=[-b-sqrt(discriminant)]/2a
=> x = [1+sqrt(61)]/10 and x = [1-sqrt(61)]/10